OpenStudy (anonymous):

please solve!! \[-4\left| 5-x \right|+4\le-16\]

OpenStudy (anonymous):



OpenStudy (anonymous):

theres your answer

OpenStudy (anonymous):

I disagree. \[5-x \ge 5 ~\text{and}~ 5-x \ge -5\]

OpenStudy (anonymous):

i got the same as rain m8 thx anyways

OpenStudy (anonymous):

You should also have \[x \ge 10\]

OpenStudy (anonymous):

yes, i noticed thank you m8